# Quantum Fisher information matrices from Rényi relative entropies

### v1

@@ 8 $\alpha$-$z$ Information matrices of parameterized thermal states
Let $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ be a parameterized family of positive definite states. A quantum generalization of the Fisher information matrix is denoted by $I^{\zeta}(\theta)$, and it has the form stated in Theorem 1, wherein the properties of the function $\zeta(x,y)$ are stated. For convenience, let us restate the form of $I^{\zeta}(\theta)$ here:
$\left[I^{\zeta}(\theta)\right]_{i,j}=\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right],$ (8.1)
@@
As stated above, Theorem 31 follows in part from Lemma 32, which establishes a formula for the information matrix $I^{\zeta}(\theta)$ of parameterized thermal states, under the assumption that a certain Fourier transform exists. Additionally, Theorem 31 follows from Lemma 33, which precisely determines the needed Fourier transform for all $\alpha\in\left(0,1\right)$ and $z>0$ when the function $\zeta(x,y)$ is set to $\zeta_{\alpha,z}\!\left(x,y\right)$, as defined in (5.6).
@@ 8.1 General formula for information matrices of parameterized thermal states
$\left[I^{\zeta}(\theta)\right]_{i,j}=\frac{1}{2}\left\langle\left\{\Phi_{f,\theta}\!\left(H_{i}\right),H_{j}\right\}\right\rangle_{\rho(\theta)}-\kappa\left\langle H_{i}\right\rangle_{\rho(\theta)}\left\langle H_{j}\right\rangle_{\rho(\theta)},$ (8.11)
where $I^{\zeta}(\theta)$ is defined in (8.1),
@@
$\left[I^{\zeta}(\theta)\right]_{i,j}$ $=\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right]$ (8.27)
@@ 8.2 $\alpha$-$z$ High-peak tent probability densities
In this subsection, I prove Lemma 33, which is the second ingredient needed for the proof of Theorem 31. The probability density in (8.73) is known as the high-peak tent, a name given in [PKPW24], due to its form when plotted (see [PKPW24, Figure 3]). The family of probability densities in (8.74)–(8.75), parameterized by $\alpha\in\left(0,1\right)$ and $z>0$, have a similar form when plotted, and so let us refer to them as the $\alpha$-$z$ high-peak tent probability densities.
@@
where (8.85) follows from [PKPW24, Lemma 12] and the function $t\mapsto\frac{2}{\pi}\ln\!\left|\coth\!\left(\frac{\pi t}{2}\right)\right|$ is a probability density function on $\mathbb{R}$ known as the high-peak tent [PKPW24]. The equality in (8.85) was calculated in [AAKS21, Section 5.1, Supplementary Information] by means of contour integration, but no argument was provided there for why the line integral along the arc of the semicircle vanishes in the large radius limit. By setting $\alpha=z=\frac{1}{2}$, the following equality holds:
@@
Setting $u=x+iy$, consider that
@@
and the fact that the function $x\mapsto\frac{\sinh\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\left(\left(\frac{\alpha}{2z}\right)x\right)}{\sinh\left(\left(\frac{1}{2z}\right)x\right)}$ is increasing on $x\in\mathbb{R}$. To see this, let us calculate the derivative of this function with respect to $x$ and prove that it is non-negative. Consider that
$\frac{\partial}{\partial x}\left(\frac{\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)}{\sinh\!\left(\left(\frac{1}{2z}\right)x\right)}\right)$
$=\frac{\left(\frac{1-\alpha}{2z}\right)\cosh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)+\left(\frac{\alpha}{2z}\right)\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\cosh\!\left(\left(\frac{\alpha}{2z}\right)x\right)}{\sinh\left(\left(\frac{1}{2z}\right)x\right)}$
$\qquad-\frac{\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)}{\left[\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\right]^{2}}\left(\frac{1}{2z}\right)\cosh\!\left(\left(\frac{1}{2z}\right)x\right)$ (8.122)
$=\frac{\left(\frac{1}{2z}\right)\cosh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)+\left(\frac{\alpha}{2z}\right)\sinh\!\left(\left(\frac{1-2\alpha}{2z}\right)x\right)}{\sinh\left(\left(\frac{1}{2z}\right)x\right)}$ (8.123)
$\qquad-\frac{\left(\frac{1}{2z}\right)\cosh\!\left(\left(\frac{1}{2z}\right)x\right)\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)}{\left[\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\right]^{2}}$ (8.124)
$=\frac{\left[\begin{array}[]{c}\left(\frac{1}{2z}\right)\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\cosh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)+\left(\frac{\alpha}{2z}\right)\sinh\!\left(\left(\frac{1-2\alpha}{2z}\right)x\right)\\ -\left(\frac{1}{2z}\right)\cosh\!\left(\left(\frac{1}{2z}\right)x\right)\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)\end{array}\right]}{\left[\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\right]^{2}}$ (8.127)
$=\frac{\left[\begin{array}[]{c}\left(\frac{1}{2z}\right)\left[\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\cosh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)-\cosh\!\left(\left(\frac{1}{2z}\right)x\right)\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\right]\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)\\ +\left(\frac{\alpha}{2z}\right)\sinh\!\left(\left(\frac{1-2\alpha}{2z}\right)x\right)\end{array}\right]}{\left[\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\right]^{2}}$ (8.130)
$=\frac{\left(\frac{1}{2z}\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)+\left(\frac{\alpha}{2z}\right)\sinh\!\left(\left(\frac{1-2\alpha}{2z}\right)x\right)}{\left[\sinh\!\left(\left(\frac{1}{2z}\right)x\right)\right]^{2}},$ (8.131)
where we made use of the identity $\sinh(x-y)=\sinh(x)\cosh(y)-\cosh(x)\sinh(y)$ twice in the above. By inspection, the last line is non-negative, implying the desired claim that $x\mapsto\frac{\sinh\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\left(\left(\frac{\alpha}{2z}\right)x\right)}{\sinh\left(\left(\frac{1}{2z}\right)x\right)}$ is increasing on $x\in\mathbb{R}$. Thus, it follows that

### v2

@@ 8 $\alpha$-$z$ Information matrices of parameterized thermal states
Let $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ be a parameterized family of positive definite states. A quantum generalization of the Fisher information matrix is denoted by $I_{\zeta}(\theta)$, and it has the form stated in Theorem 1, wherein the properties of the function $\zeta(x,y)$ are stated. For convenience, let us restate the form of $I_{\zeta}(\theta)$ here:
$\left[I_{\zeta}(\theta)\right]_{i,j}=\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right],$ (8.1)
@@
As stated above, Theorem 32 follows in part from Lemma 33, which establishes a formula for the information matrix $I_{\zeta}(\theta)$ of parameterized thermal states, under the assumption that a certain Fourier transform exists. Additionally, Theorem 32 follows from Lemma 34, which precisely determines the needed Fourier transform for all $\alpha\in\left(0,1\right)$ and $z>0$ when the function $\zeta(x,y)$ is set to $\zeta_{\alpha,z}\!\left(x,y\right)$, as defined in (5.6).
@@ 8.1 General formula for information matrices of parameterized thermal states
$\left[I_{\zeta}(\theta)\right]_{i,j}=\frac{1}{2}\left\langle\left\{\Phi_{f,\theta}\!\left(H_{i}\right),H_{j}\right\}\right\rangle_{\rho(\theta)}-\kappa\left\langle H_{i}\right\rangle_{\rho(\theta)}\left\langle H_{j}\right\rangle_{\rho(\theta)},$ (8.11)
where $I_{\zeta}(\theta)$ is defined in (8.1),
@@
$\left[I_{\zeta}(\theta)\right]_{i,j}$ $=\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right]$ (8.27)
@@ 8.2 $\alpha$-$z$ High-peak tent probability densities
In this subsection, I prove Lemma 34, which is the second ingredient needed for the proof of Theorem 32. The probability density in (8.73) is known as the high-peak tent, a name given in [PKPW24], due to its form when plotted (see [PKPW24, Figure 3]). It was shown in [EO19, Appendix C], [AAKS21, Section 5.1, Supplementary Information], and [PKPW24, Lemma 12] that the function $p(t)$ is the Fourier transform of $\tanh\!\left(\frac{\omega}{2}\right)/\left(\frac{\omega}{2}\right)$, the latter function having been considered in the context of quantum belief propagation [Has07, Kim12, KB19].
The family of probability densities in (8.74)–(8.75), parameterized by $\alpha\in\left(0,1\right)$ and $z>0$, have a similar form to $p(t)$ when plotted, and so let us refer to them as the $\alpha$-$z$ high-peak tent probability densities.
@@
  3. 3.
    Similarly, for all $\alpha\in\left(0,1\right)$, the characteristic function of $p_{\alpha,z}(t)$ converges pointwise to one everywhere in the limit $z\to\infty$, implying that, for all $\alpha\in\left(0,1\right)$, the random variable $X_{\alpha,z}\sim q_{\alpha,z}$ converges in distribution to the random variable $X\sim p$ in the limit $z\to\infty$.
@@
where (8.85) follows from [PKPW24, Lemma 12] and the function $t\mapsto\frac{2}{\pi}\ln\!\left|\coth\!\left(\frac{\pi t}{2}\right)\right|$ is a probability density function on $\mathbb{R}$ known as the high-peak tent [PKPW24]. The equality in (8.85) was calculated in [EO19, Appendix C] and [AAKS21, Section 5.1, Supplementary Information] by means of contour integration. By setting $\alpha=z=\frac{1}{2}$, the following equality holds:
@@
To see (8.100), setting $u=x+iy$, consider that
@@
and the fact that the function $x\mapsto\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)/\sinh\!\left(\left(\frac{1}{2z}\right)x\right)$ is increasing on $x\in\mathbb{R}$. Indeed, since this latter function is odd, to see that it is increasing on $x\in\mathbb{R}$, it suffices to prove that it is increasing on $\left(0,\infty\right)$. To see this, let us calculate the derivative of this function with respect to $x$ and prove that it is non-negative on $\left(0,\infty\right)$. Consider that
$2\frac{\partial}{\partial x}\left(\frac{\sinh\!\left(\left(\frac{1-\alpha}{2z}\right)x\right)\sinh\!\left(\left(\frac{\alpha}{2z}\right)x\right)}{\sinh\!\left(\left(\frac{1}{2z}\right)x\right)}\right)$
$=\frac{\partial}{\partial x}\left(\frac{\left(e^{\left(\frac{1-\alpha}{2z}\right)x}-e^{-\left(\frac{1-\alpha}{2z}\right)x}\right)\left(e^{\frac{\alpha}{2z}x}-e^{-\frac{\alpha}{2z}x}\right)}{e^{\frac{1}{2z}x}-e^{-\frac{1}{2z}x}}\right)$ (8.122)
$=\frac{\partial}{\partial x}\left(\frac{\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{e^{\frac{1}{z}x}-1}\right)$ (8.123)
$=\frac{\frac{1-\alpha}{z}e^{\left(\frac{1-\alpha}{z}\right)x}\left(e^{\frac{\alpha}{z}x}-1\right)+\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)e^{\frac{\alpha}{z}x}}{e^{\frac{1}{z}x}-1}-\frac{\frac{1}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)e^{\frac{1}{z}x}}{\left(e^{\frac{1}{z}x}-1\right)^{2}}$ (8.124)
$=\frac{\frac{1-\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1+1\right)\left(e^{\frac{\alpha}{z}x}-1\right)+\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1+1\right)}{e^{\frac{1}{z}x}-1}$
$\qquad-\frac{\frac{1}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)e^{\frac{1}{z}x}}{\left(e^{\frac{1}{z}x}-1\right)^{2}}$ (8.125)
$=\frac{\left[\begin{array}[]{c}\frac{1-\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)+\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)\\ +\frac{1-\alpha}{z}\left(e^{\frac{\alpha}{z}x}-1\right)+\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\end{array}\right]}{e^{\frac{1}{z}x}-1}$ (8.128)
$\qquad-\frac{\frac{1}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{\left(e^{\frac{1}{z}x}-1\right)}\frac{e^{\frac{1}{z}x}}{\left(e^{\frac{1}{z}x}-1\right)}$ (8.129)
$=\frac{\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{\left(e^{\frac{1}{z}x}-1\right)}\left(\frac{1-\alpha}{z}+\frac{\alpha}{z}-\frac{e^{\frac{1}{z}x}}{z\left(e^{\frac{1}{z}x}-1\right)}\right)$
$\qquad+\frac{\frac{1-\alpha}{z}\left(e^{\frac{\alpha}{z}x}-1\right)+\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)}{e^{\frac{1}{z}x}-1}$ (8.130)
$=\frac{\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{\left(e^{\frac{1}{z}x}-1\right)}\frac{1}{z}\left(1-\frac{e^{\frac{1}{z}x}}{e^{\frac{1}{z}x}-1}\right)+\frac{\frac{1-\alpha}{z}\left(e^{\frac{\alpha}{z}x}-1\right)+\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)}{e^{\frac{1}{z}x}-1}$ (8.131)
$=-\frac{\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{\left(e^{\frac{1}{z}x}-1\right)}\left(\frac{1}{z\left(e^{\frac{1}{z}x}-1\right)}\right)+\frac{\frac{1-\alpha}{z}\left(e^{\frac{\alpha}{z}x}-1\right)}{e^{\frac{1}{z}x}-1}+\frac{\frac{\alpha}{z}\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)}{e^{\frac{1}{z}x}-1}$ (8.132)
$=\frac{\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{z\left(e^{\frac{1}{z}x}-1\right)}\left(\frac{1-\alpha}{e^{\left(\frac{1-\alpha}{z}\right)x}-1}+\frac{\alpha}{e^{\frac{\alpha}{z}x}-1}-\frac{1}{e^{\frac{1}{z}x}-1}\right).$ (8.133)
Since $\frac{\left(e^{\left(\frac{1-\alpha}{z}\right)x}-1\right)\left(e^{\frac{\alpha}{z}x}-1\right)}{z\left(e^{\frac{1}{z}x}-1\right)}>0$ for all $x>0$, the desired statement will follow if the inequality
$\frac{1-\alpha}{e^{\left(\frac{1-\alpha}{z}\right)x}-1}+\frac{\alpha}{e^{\frac{\alpha}{z}x}-1}\geq\frac{1}{e^{\frac{1}{z}x}-1}$ (8.134)
holds for all $\alpha\in\left(0,1\right)$ and $x>0$. Thus, the statement follows if the function $x\mapsto\frac{1}{e^{\frac{1}{z}x}-1}$ is convex and decreasing on $x\in\left(0,\infty\right)$. Indeed, if these properties hold, then
$\frac{1-\alpha}{e^{\left(\frac{1-\alpha}{z}\right)x}-1}+\frac{\alpha}{e^{\frac{\alpha}{z}x}-1}\geq\frac{1}{e^{\frac{1}{z}\left(\left(1-\alpha\right)^{2}+\alpha^{2}\right)x}-1}\geq\frac{1}{e^{\frac{1}{z}x}-1},$ (8.135)
where the first inequality follows from convexity and the second from the decreasing property because $\left(1-\alpha\right)^{2}+\alpha^{2}<1$ for $\alpha\in\left(0,1\right)$. To prove these properties, let us calculate the first and second derivatives of the function $x\mapsto\frac{1}{e^{\frac{1}{z}x}-1}$:
$\frac{\partial}{\partial x}\left(\frac{1}{e^{\frac{1}{z}x}-1}\right)$ $=-\frac{e^{\frac{x}{z}}}{z\left(e^{\frac{1}{z}x}-1\right)^{2}},$ (8.136)
$\frac{\partial^{2}}{\partial x^{2}}\left(\frac{1}{e^{\frac{1}{z}x}-1}\right)$ $=\frac{e^{\frac{x}{z}}\left(e^{\frac{x}{z}}+1\right)}{z^{2}\left(e^{\frac{1}{z}x}-1\right)^{3}}.$ (8.137)
Since $\frac{\partial}{\partial x}\left(\frac{1}{e^{\frac{1}{z}x}-1}\right)\leq 0$ and $\frac{\partial^{2}}{\partial x^{2}}\left(\frac{1}{e^{\frac{1}{z}x}-1}\right)\geq 0$ on $x\in\left(0,\infty\right)$, the claim follows. Thus, it follows that